OpenStudy (anonymous):

show the differentiation rule for: (cosx)' = -sinx for any x answer asap please @Mathematics

OpenStudy (anonymous):

lim(h-->0) (cos(x+h) - cosx)/h = lim(h-->0) (cosxcosh - sinxsinh - cosx)/h = lim(h-->0) ((cosxcosh-cosx)/h - sinxsinh/h) = lim(h-->0) cosx(cosh-1/h) - sinx(sinh/h)) = cosx(0) - sinx(1) = -sinx

OpenStudy (anonymous):

cosh = 1 when h = 0 1-1 = 0 0/0 =0 Also, sinh/h = 1

OpenStudy (anonymous):

when h = 0